OpenStudy (rachie19):

64^2x-3=1/32^5. solve for x...please help

OpenStudy (rachie19):

\[64^{2x-3}= \frac{ 1 }{ 32^{5} }\]

hartnn (hartnn):

what have you tried?

OpenStudy (rachie19):

i used 2 so that i could have the same base but its not workinh

hartnn (hartnn):

2 is the correct base :) 64 = 2^6 32 = 2^5 Also we will need \(\dfrac{1}{x^n} = x^{-n}\) it would be helpful if you show your work :)

OpenStudy (rachie19):

i think im more stuck on how to write the actual problem

OpenStudy (rachie19):

6(2x-3)=?

hartnn (hartnn):

\(64 =2^6 \\ \Large 64^{2x-3} = (2^6)^{2x-3} = 2^{6(2x-3)}\)

hartnn (hartnn):

\(\dfrac{1}{32^5 } = 32^{-5} = (2^5)^{-5} = 2^{5\times(-5)}= ...\)

hartnn (hartnn):

so, comparing the exponents 6(2x-3) = 5*(-5)

OpenStudy (rachie19):

thank you

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

let me know if you have any more doubts in this problem :)

OpenStudy (rachie19):

i got -13/7 is that what you got too?

OpenStudy (rachie19):

@hartnn

hartnn (hartnn):

12x-18 = -25 12x = 18-25 12x = -7 x = -7/12 how did you get -13/7 ?

OpenStudy (rachie19):

at first i got 19/4 then i got -13/7

hartnn (hartnn):

if you show your work, i can help you to spot the error :)

OpenStudy (rachie19):

ok i see what i did wrong.thanks you again =)

hartnn (hartnn):

welcome ^_^